# Understanding the Dataset Practitioners Behind Large Language Model Development

### v1

@@ Abstract.
As large language models (LLMs) become more advanced and impactful, it is increasingly important to scrutinize the data that they rely upon and produce. What is it to be a dataset practitioner doing this work? We approach this in two parts: first, we define the role of ”dataset practitioner” by performing a retrospective analysis on the responsibilities of teams contributing to LLM development at Google. Then, we conduct semi-structured interviews with a cross-section of these practitioners (N=10). We find that data quality is the top priority. To evaluate data quality, practitioners either rely on their own intuition or write custom evaluation logic. There is a lack of consensus across practitioners on what quality is and how to evaluate it. We discuss potential reasons for this phenomenon and opportunities for alignment.
@@ 1. Introduction
As the state-of-the-art for large language models (LLMs) advances (OpenAI, 2023), the field of relevant data analysis is rapidly evolving. Because the data used and produced by LLMs is largely unstructured, traditional statistical analyses are insufficient for rigorous evaluation (Wach et al., 2023; Sambasivan et al., 2021; Brown et al., 2020). Furthermore, as applications of these LLMs become more widely adopted and impactful (Abdullah et al., 2022; Baldassarre et al., 2023), there’s a deeper need to qualitatively understand these datasets; for instance, to mitigate sociological biases, ensure safe outputs, and minimize harm.
We aim to identify the needs and challenges of those who want to understand unstructured, text-based datasets for LLM development: a group that we define as dataset practitioners. To develop this definition, we perform a retrospective analysis within Google. We then conduct semi-structured interviews with a cross-section of these practitioners (N=10) to better understand their workflows, tools, and challenges.
We find that practitioners prioritize data quality; however, there is no consensus on what constitutes ”high quality” data. Despite HCI/VIS researchers’ active efforts to deliver relevant sensemaking methods and tools, data practitioners in aggregate do not appear to be adopting these solutions, instead relying either on cursory visual inspection of spreadsheets or custom analyses logic in notebooks to understand their data. There is demand for frameworks, consensus, and tooling in this space that is not being met. We discuss hypotheses for this observed phenomenon, and conclude with opportunities for further research and alignment.
@@ 2.1. Techniques and tools
To better understand unstructured text outputs of LLMs, there is work in developing new explainability algorithms and techniques (Das and Rad, 2020; Danilevsky et al., 2020; Kim et al., 2018). These techniques can be packaged in frameworks and tools (Kandel et al., 2012; Agarwal and Das, 2020) that practitioners can use to explore data with varying degrees of effort. Low-effort exploration involves cursory, surface-level viewing of data, with minimal interaction or manipulation. These interactions typically occur in tools used to view tabular data, such as Google Sheets or Microsoft Excel. On the other end, high-effort exploration may involve writing code to perform custom analyses. This is commonly done in Python scripts or notebooks such as Google Colab or Jupyter (Chattopadhyay et al., 2020; Kery et al., 2019; Tabard et al., 2008; Kery et al., 2018). We define intermediate-effort exploration as involving more effort than opening up a spreadsheet, but less effort than writing custom code. This category might involve actions such as interacting with a user interface or running existing scripts. Using more interactive features in Google Sheets, such as developing formulas or annotating data, would fall in this category. There are a trove of curated tools for data understanding and model interpretability at this level of exploration, including the Language Interpretability Tool (Tenney et al., 2020) and What-If Tool (Wexler et al., 2020), Model Tracker (Amershi et al., 2015), ChainForge (Arawjo et al., 2023), RuleMatrix (Bellamy et al., 2018), and AI Fairness 360 (Ming et al., 2019).
2.2. User studies
User studies on practitioner-tooling interaction have identified cognitive biases in the workflow such as selective exploration of data, over-trust and misuse of interpretability tools, and gaps in understanding regarding data visualizations (Kaur et al., 2020; Doshi-Velez and Kim, 2017; Miller, 2019).
2.3. Data priorities
Datasets relevant to LLM development have become increasingly composed of smaller, curated subsets that target address specific concerns, such as safety and fairness (Mehrabi et al., 2021). The focus is increasingly on data quality (Sambasivan et al., 2021) rather than quantity (Durrant-Whyte, 2015); though quantifying the criteria for quality is an open problem (Herman, 2019; Doshi-Velez and Kim, 2017).
2.4. Analyzing analyzers
Previous work has classified those who work with data as data workers or data scientists (Harris et al., 2013; Wang et al., 2019; Han et al., 2020; Crisan et al., 2021; Muller et al., 2019). Data worker satisfies the breadth that aim to capture, as our target population spans more job titles than data scientist. Data scientist is too narrow for our population. It does not encompass the specific challenges introduced by the new LLM-centered data regime, such as a rising need for qualitative evaluation methods or the broader range of job responsibilities within this role. These broader responsibilities might include, for example, creating new architecture to interpret data, or developing adjudication methods for human-labelled data.
@@ 3. Retrospective analysis
To identify the data practitioner, we conducted a retrospective analysis of teams working on developing LLMs within Google. Google’s organizational structure is uniquely positioned to support a broad survey of the landscape as the technology stack is vertically-integrated (Harrigan, 1985); that is, the relevant tooling, infrastructure, modeling, evaluation, and research are primary developed in-house. Due to the technology stack’s vertical integration, we were able to mitigate selection bias in our sample by maximizing exposure to a diverse range of projects.
Using company-internal organizational charts and employee directories, we identified projects associated with the development of Google’s core LLMs. We also conducted a meta-review of company-internal user studies around evaluating tools for data exploration. Applying a grounded theory methodology (Corbin and Strauss, 1990), we inductively applied a relational content analysis and synthesized common themes to develop a framework around data practitioning.
@@ 3.1. Defining the dataset practitioner
Our dataset practitioner seeks to understand unstructured, text-based data for the purpose of developing large language models. The day-to-day work of the practitioner can cover a broad range of tasks traditionally defined in roles such as software developer, machine learning engineer, data scientist, research scientist, product manager, and product counsel. The practitioner may prioritize these responsibilities concurrently, or switch gears along the model development lifecycle. They may do any of the following representative tasks:
@@
Analyzing a dataset (user feedback, comments, etc) to find trends in it
@@
Identifying key metrics to define “quality” for a specific use case
Next, we given examples of datasets that they may explore. The term ”dataset” can traditionally imply a static and well-curated set of data; we expand this notion to include any set of text examples, which may come from a variety of provenances (e.g. scraped, synthetically generated, curated by experts). They fall in to two broad categories:
@@
Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) data: These datasets are significantly smaller and more specialized than pre-training data, and are used to refine the LLM from an open-ended generation model to a specific use case- most notably, the chatbot interface that many productionized LLMs employ. LLMs can be fine-tuned for other specific products and use cases.
@@
Model outputs: Model outputs can be evaluated outside of the context of a specific benchmark. Side-by-side analysis of model outputs may be conducted against golden sets or outputs from a baseline model.
@@ 4.1. Participants
Using our updated definition, we recruited ten dataset practitioners (N=10) across different domains of the model development lifecycle (Crisan et al., 2021; Kandel et al., 2012); broadly: tooling, modeling, and evaluation. These participants and their current focus areas are listed in Table 1. We validated our observation from Section 3.1 that the domains of their work are fluid; participants who identified in one domain during our recruiting cycle demonstrated experience in many adjacent areas within the interview. For example, a practitioner formerly focused on modeling shifted priorities to safety and fairness evaluation as their models became more scrutinized and regulated, and two tool-builders reported being driven to build tooling to address their own unmet needs in modeling.
[…]
@@
4.2. Interviews
Following recruitment and an informed consent process, we conducted semi-structured, one-on-one interviews with participants over video conferencing. Each 30-minute interview covered three broad topics:
@@ 4.2. Interviews
Tools and techniques: Awareness and usage of existing tools and pipelines, decision making,
advantages and limitations, statistical and visual interpretability methods -
@@
We synthesized our findings through a thematic analysis (Braun and Clarke, 2006). Each interview was de-identified, transcribed, broken into excerpts, and coded. The interview subject matter and codebook were adopted from previous informal contextual inquiries and interview protocols from adjacent research studies (Kaur et al., 2020; Li et al., 2023). Thematic elements, behaviors, and representative quotes in this paper are saturated (Guest et al., 2006; Ando et al., 2014), with a code repeated in at least three distinct transcriptions.
5. Results
5.1. Assessing data quality is the biggest challenge.
There has been a recent shift in the literature from prioritizing benchmarks and abstract metrics (Wagstaff, 2012) to prioritizing data quality (Sambasivan et al., 2021). We observed this phenomenon in our sample as well; data quality—defining, finding, and identifying high-quality data—was unanimously the biggest user challenge and priority across all use cases (Table 2, Challenges).
@@ 5. Results
5.2. However, practitioners rely largely on their own intuition to validate this data quality.
Despite this focus on high-quality data, all participants mentioned understanding their datasets by viewing it in tabular form in spreadsheets (Section 2.1, low-effort exploration); that is, they’re looking at a handful of examples.
@@
Participants cited efficiency, customization, a short learning curve, and ease-of-sharing as reasons why they rely on spreadsheets (Table 2, Challenges).111Interestingly and consistent with similar user studies, our participants emphasized that their reliance on visual inspection of spreadsheets were their own behaviors and not best practices. They suggested that other practitioners likely used more sophisticated tooling (Pennebaker, 2011).
5.3. Or, practitioners will run custom analyses.
Participants were not engaging in low-effort exploration for lack of technical know-how; 7 of the 9 participants mentioned writing custom code in Python notebooks to explore their data (Section 2.1, high-effort exploration), and in one instance even to train production models. Participants liked the ease of sharing and customizing these notebooks, and cited reliability, setup, code management as pain points of these notebooks (Table 2, Challenges), validating results from other studies on Python notebook usage (Kery et al., 2018; Chattopadhyay et al., 2020; Kery et al., 2019; Tabard et al., 2008).
5.4. Practitioners recognize the confirmation biases in their exploration practices.
The majority— if not all— of the data exploration is being done between visual inspection in spreadsheets and custom logic in Python notebooks, These two practices have been observed to perpetuate implicit cognitive bias in other settings (Herman, 2019; Pirolli and Card, 2005; Gilpin et al., 2018; Caliskan et al., 2017). Indeed, our participants admit to this confirmation bias in their practices:
““I look at data …eyeballing that things make sense.” —M2
Model developers reported that they did not look at training data unless their model outputs were surprising.
@@
5.5. There is missing alignment on intermediate-effort exploration patterns.
Practitioners across different domains have converged upon a low-effort exploration pattern: visually scanning data in Google Sheets. They’ve also converged upon a high-effort exploration pattern: writing custom Python code in Colab notebooks. However, there were no tools and processes within our description of (Section 2.1, intermediate-effort exploration) that achieved inter-practitioner alignment. The few practitioners who performed intermediate-effort exploration did so using different methods, such as running a binary for calculating safety and toxicity thresholds, kicking off a pipeline to automatically classify their data, and using a UI to visualize embeddings.
@@
This lack of alignment poses an organizational challenge; as mentioned in Section 2.3, LLM training datasets are increasingly composed of smaller datsets to leverage the expertise of specific subteams.
@@ 6. Discussion
The reason why practitioners have not aligned on an intermediate-level exploration pattern is not obvious. Practitioners across all domains recognize that there is a gap in the workflow:
@@
“Right now, if you want to curate high-quality data, you go through [each point] manually as an expert, which is not scalable on the scale of thousands of examples.” —T2
They are aware of and have tried the existing tools in this space. They are aligned on the properties that they want out of this tool (Table 2, Challenges), and these requests are being communicated to tooling teams:
“The kinds of requests we tend to get nowadays are about larger-scale dataset management, like mixture building. When you have a big selection, going through and reviewing 10,000 rows is not what you want to do …That is much more amenable to summary review.” —T1
@@
6.1. The toolmakers’ hypothesis: the world is new.
When tool developers (T1-T4) described exploration workflows, they explained that there was a lack of alignment because the field is still taking shape:
@@
Second, our participants described a landscape where there was a lack of alignment across multiple topics such as objectives, metrics, and benchmarks. This is validated in the literature (Doshi-Velez and Kim, 2017; Gilpin et al., 2018), and suggest that the field and its principles are still emerging.
@@
“There’s so many competing definitions of prompt quality…it’s a research north star that happens to be a major product priority. How can we improve this extremely important data set?’ —T1, on LLM prompts
“The quality of data is subjective; a lot of people disagree…one person thinks it’s really high-quality data, but there’s no objective.” —M1, on training data
“There’s not a framework for evaluating [training data].. in a perfect world, there is well-articulated behavior (tone, subject matter, objective results)..” —T3, on evaluation data
@@
“[Consider] search rankings…what makes for a good benchmark, how do we come to an agreement?” —M1
“If you’re doing simple classification, it’s easy to measure accuracy or precision or recall. But with these generative models, evaluation is very subjective. Even the output of the model is subjective, so then, what’s going into the model- it’s really hard to say, is this better or worse?” —E1
@@
“Think about safety data curation.. people can’t agree on criteria, let alone apply that criteria at scale.” —T2
@@
“What [data practitioners are] actually doing and what they communicate that they need are two very different things. What are they actually trying to do?” —T2
This lack of alignment is a bigger problem now that teams are collaborating more closely to curate training datasets. Even if one team in the development pipeline identifies their quality evaluation parameters, there needs to be further agreement at the inter-team level.
6.2. The model developers’ hypothesis: there’s no tool that works for my use case.
Practitioners in the modeling and evaluation domains speculated that alignment was unlikely due to the custom needs and requirements across use cases (Section 3.1).
“I think why [a spreadsheet is] so universal is that it’s so basic.. you can customize it yourself to give this affordance that other tools may not give you.. it’s simple.” —E1
“We have tried so many [tools]. Why I think these tools are limiting is because they offer you exploration on only one aspect of [the data]…For me, they’re too specific.” —M2
Interestingly, when asked about the custom requirements specific to their use cases, practitioners listed many similar requirements, including:
@@
It is likely the case that both hypotheses are true to some extent. There may be select opportunities for alignment as the field matures, and there are likely other problems that will require custom solutions. For example, there are specific tools being developed to address challenges that persist across datasets, such as safety and toxicity classification (Bellamy et al., 2018).
7. Future work and conclusions

### v2

@@ Abstract.
As large language models (LLMs) become more advanced and impactful, it is increasingly important to scrutinize the data that they rely upon and produce. What is it to be a dataset practitioner doing this work? We approach this in two parts: first, we define the role of “dataset practitioners” by performing a retrospective analysis on the responsibilities of teams contributing to LLM development at a technology company, Google. Then, we conduct semi-structured interviews with a cross-section of these practitioners (N=10). We find that although data quality is a top priority, there is little consensus around what data quality is and how to evaluate it. Consequently, practitioners either rely on their own intuition or write custom code to evaluate their data. We discuss potential reasons for this phenomenon and opportunities for alignment.
@@ 1. Introduction
As the state-of-the-art for large language models (LLMs) advances (OpenAI, 2023; Team et al., 2023), the field of relevant data analysis is rapidly evolving. Because the data used and produced by LLMs is largely unstructured, traditional statistical analyses are insufficient for rigorous evaluation (Wach et al., 2023; Sambasivan et al., 2021; Brown et al., 2020). Furthermore, as applications of these LLMs become more widely adopted and impactful (Abdullah et al., 2022; Baldassarre et al., 2023), there is a deeper need to qualitatively understand these datasets; for instance, to mitigate sociological biases, ensure safe outputs, and minimize harm.
We aim to identify the needs and challenges of those who want to understand unstructured, text-based datasets for LLM development: a group that we define as dataset practitioners. To develop this definition, we perform a retrospective analysis within Google, a technology company that is developing LLMs. We then conduct semi-structured interviews with a cross-section of practitioners (N=10) to better understand their workflows, tools, and challenges.
We find that practitioners increasingly prioritize data quality; however, there is no consensus on what constitutes “high quality” data. Despite the HCI and visualization researchers’ active efforts to deliver relevant sensemaking methods and tools, data practitioners in aggregate do not appear to be adopting these solutions, instead relying either on cursory visual inspection of spreadsheets or custom analyses logic in notebooks to understand their data. There is demand for frameworks, consensus, and tooling in this space that is not being met. We discuss hypotheses for this observed phenomenon, and conclude with opportunities for further research and alignment.
@@ 2. Related Work
2.1. Analyzing Analyzers
As data science has grown as a discipline, so have the amount of analyses (Harris et al., 2013; Crisan et al., 2021), surveys (Zhang et al., 2020), and interviews (Wang et al., 2019) performed to capture the role of those who do this work.
Some notable highlights include Kandel et al. (Kandel et al., 2012), which classifies the emerging role of the data analysts across different industrial sectors, such as healthcare and retail. Muller et al. (Muller et al., 2019) interviewed data scientists at IBM to capture different approaches to their work, and Crisan et al. (Crisan et al., 2021) creates a taxonomy of job roles across data workers, such as moonlighters, generalists or evangelists.
Across these studies, the definitions of data analysts or data workers satisfy the breadth of work that we aim to capture in this inquiry. Data scientist is too narrow for our population. It does not encompass the specific challenges introduced by the new LLM-centered data regime, such as a rising need for qualitative evaluation methods or the broader range of job responsibilities within this role. These broader responsibilities might include, for example, creating new architecture to interpret data, or developing adjudication methods for human-labelled data.
@@ 2.2. Techniques and Tools
There have also been existing inquiries into the techniques and tools that practitioners use. Many data science workers interact with data in tabular formats, using tools such as Google Sheets or Microsoft Excel (Birch et al., 2018). They may also writing code to perform custom analyses, commonly by using Python scripts or notebooks such as Google Colab or Jupyter (Chattopadhyay et al., 2020; Kery et al., 2019; Tabard et al., 2008; Kery et al., 2018).
As large language models have become more salient, the space of applicable techniques and tools has increased. The field of explainable AI (XAI) (Das and Rad, 2020; Danilevsky et al., 2020) has yielded new explainability (Kim et al., 2018; Ribeiro et al., 2016) and visualization techniques for natural language processing. These techniques can be packaged into frameworks and tools (Kandel et al., 2012; Agarwal and Das, 2020; Ashtari et al., 2023), such as Language Interpretability Tool (Tenney et al., 2020), What-If Tool (Wexler et al., 2020), and AI Fairness 360 (Bellamy et al., 2018) among many others (Amershi et al., 2015; Arawjo et al., 2023; Ming et al., 2019; Hohman et al., 2018; La Rosa et al., 2023). However, these LLM-focused tools are relatively recent, and there is a lack of existing research assessing the extent of their adoption across industry and academia.
2.3. Curation Trends
Datasets relevant to LLM development have become increasingly composed of smaller, curated subsets that target address specific concerns, such as safety and fairness (Mehrabi et al., 2021; Wagstaff, 2012). The focus is increasingly on data quality (Sambasivan et al., 2021) rather than quantity (Durrant-Whyte, 2015), though quantifying the criteria for data quality is an open problem (Herman, 2019; Doshi-Velez and Kim, 2017).
@@ 3. Retrospective analysis
To define the role of data practitioner, we conducted a retrospective analysis of teams working on developing LLMs at Google. This company’s organizational structure is uniquely positioned to support a broad survey of the landscape as the technology stack is vertically-integrated (Harrigan, 1985); that is, the relevant tooling, infrastructure, modeling, evaluation, and research are primary developed in-house. For example, Google has infrastructure teams that build custom software to deploy ML experiments on computational resources, tooling teams that create applications for interpreting model outputs, data teams that source and clean human data, modeling teams that improve LLM models across different modalities, and safety teams that focus on enforcing policies and model quality.
Using company-internal organizational charts and employee directories, we identified projects associated with the development of the company’s core LLMs. We also conducted a meta-review of company-internal user studies around evaluating tools for data exploration. Applying a grounded theory methodology (Corbin and Strauss, 1990), we inductively applied a relational content analysis and synthesized common themes to develop a framework around data practitioning.
@@ 3.1. Defining the Dataset Practitioner
The dataset practitioner interacts with unstructured, text-based data for the purpose of developing large language models. The practitioner’s day-to-day work can cover a broad range of tasks traditionally defined in roles such as software developer, machine learning engineer, data scientist, research scientist, product manager, or product counsel. The practitioner may prioritize these responsibilities concurrently, or switch gears along the model development lifecycle. They may do any of the following representative tasks:
@@
Analyzing a dataset (feedback, comments, etc) to find trends
@@
Identifying key metrics to define “quality” for a use case
Next, we give examples of datasets that they may explore. The term “dataset” traditionally implies static and well-curated data; we expand this notion to include any set of text examples, which may come from a variety of provenances (e.g. scraped, synthetically generated, curated by experts). We categorize these broadly:
@@
SFT and RLHF data: Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) datasets are used to refine pre-trained LLMs (Team et al., 2023; OpenAI, 2023). They are significantly smaller and more specialized than pre-training data, and can be used from an open-ended generation model to a specific use case—most notably, the chatbot interface that many productionized LLMs employ. LLMs can be fine-tuned for other specific products and use cases as well.
@@
Model outputs: Model outputs can be evaluated outside of the context of a specific benchmark. Side-by-side analysis of model outputs may be conducted against golden sets or outputs from a baseline model (Kahng et al., 2024).
Outputs of in-context learning: These are a specific subset of model outputs. In-context learning has allowed users to create new models with no golden data at all. These may then be evaluated by analyzing the outputs from multiple runs of a prompt.
@@ 4.1. Participants
Using our updated definition, we recruited ten dataset practitioners (N=10) within Google for our study.111Note that Reif et al. (Reif et al., 2024) uses the same participant sample. We selected these participants with the criteria that their current work involves interacting with datasets for the purposes of developing large language models, and prioritized sampling participants from a variety of concentrations and backgrounds. These participants and their primary focus areas (tooling, modeling, or evaluation) are listed in Table 1. We validated our observation from Section 3.1 that the domains of their work are fluid; participants who identified in one domain during our recruiting cycle demonstrated experience in many adjacent areas within the interview. For example, a practitioner formerly focused on modeling shifted priorities to safety and fairness evaluation as their models became more scrutinized and regulated, and two tool-builders reported being driven to build tooling to address their own unmet needs in modeling.
Domain | Participant ID | Focus Area |
@@
4.2. Interview Protocol
Following recruitment and an informed consent process, we conducted semi-structured, one-on-one interviews with participants over video conferencing. Each 30-minute interview covered the following topics:
@@ 4.2. Interview Protocol
Tools and techniques: Awareness and usage of existing tools and pipelines, decision making, advantages and limitations, statistical and visual interpretability methods
@@
We curated the interview topics from prior contextual inquiries and protocols from similar research studies in defining data work (Kaur et al., 2020; Li et al., 2023; Wang et al., 2019). By following a similar interview protocol, we hope to isolate the specific challenges faced in LLM-development.
We synthesized our findings through a thematic analysis (Braun and Clarke, 2006). Each interview was de-identified, transcribed, broken into excerpts, and coded. Thematic elements, behaviors, and representative quotes in this paper are saturated (Guest et al., 2006; Ando et al., 2014), with a code repeated in at least three distinct transcriptions.
[A matrix categorizing findings.]There are 10 columns, one for each participant. Each row is a specific theme that was discovered and coded. All participants mentioned interacting with spreadsheets and cited data quality as a challenge in their work.
4.3. Findings
4.3.1. Participants prioritize data quality.
Corroborating the prior work described in Section 2.3, we find that data quality—defining, finding, and identifying high-quality data—was unanimously the biggest user challenge and priority across all use cases (Table 2, Challenges).
@@ 4.3. Findings
Although data quality has always remained an important priority for data scientists, these concerns were addressable through tasks such as data cleaning (Muller et al., 2019) or feature engineering (Crisan et al., 2021). In the context of generative modeling, the evaluation metrics and consensus frameworks are less straightforward.
4.3.2. However, practitioners rely largely on their own intuition to validate this data quality.
All participants reported that they would evaluate their data by scanning it visually in spreadsheet form; that is, they would look at a handful of examples.
@@
Participants cited efficiency, customization, a short learning curve, and ease-of-sharing as reasons for their reliance on spreadsheets (Table 2, Challenges).222Interestingly and consistent with similar user studies, our participants emphasized that their reliance on visual inspection of spreadsheets were their own behaviors and not best practices. They suggested that other practitioners likely used more sophisticated tooling (Pennebaker, 2011). While these factors align with prior research on spreadsheet usage (Birch et al., 2018), the ease-of-sharing factor may particularly encourage practitioners to use spreadsheets for LLM development. Unlike the data analysts in Kandel et al. (Kandel et al., 2012), who collaborated with “hacker”-types with scripting and coding proficiency, our participants reported needing to share data with a larger and more diverse set of stakeholders, such as directors and legal teams, to review high-stakes safety fine-tuning datasets.
4.3.3. Or, practitioners will run custom analyses.
Seven of the nine participants mentioned also writing custom code in Python notebooks to explore their data, and in one instance even to train production models. Participants liked the customization of these notebooks (Kery et al., 2019), yet cited reliability, setup, efficiency, code management as pain points (Table 2, Challenges), validating results from other studies on Python notebook usage (Kery et al., 2018; Chattopadhyay et al., 2020; Kery et al., 2019; Tabard et al., 2008).
The efficiency concerns around long-running computations in Python notebooks (Chattopadhyay et al., 2020) may be further exacerbated as LLMs require more computational power; participants mentioned that “getting model servers up and running takes forever” (R1), “my queries [to LLM APIs] take a while” (E1), and they wished they had “infinite QPS (Queries Per Second) [for their LLM API]” (R2).
4.3.4. Practitioners recognize the confirmation biases in their exploration practices.
The majority—if not all—of the data exploration is being done between visual inspection in spreadsheets and custom logic in Python notebooks, allowing the practitioner to look at whatever they would like. This degree of freedom exacerbates cognitive bias (Herman, 2019; Pirolli and Card, 2005; Gilpin et al., 2018; Caliskan et al., 2017); for example, Miller et al. (Miller, 2019) mentions that “explainable AI uses only the researchers’ intuition of what constitutes a ‘good’ explanation.” Indeed, our participants admit to this confirmation bias in their practices:
““I eyeball that things make sense [in the data].” —M2
In fact, model developers reported that they did not look at training data unless their model outputs were surprising.
@@
4.3.5. Participants have not converged upon other tools.
Apart from Google Sheets and Python notebooks like Colab, no other tools garnered consensus among practitioners. Some practitioners employed additional methods, such as running a binary for calculating safety and toxicity thresholds, kicking off a pipeline to automatically classify their data, and using an user interface to visualize embeddings. However, these practices were not prevalent in our sample.
@@
The lack of alignment in tooling presents an organization challenge. As training datasets are increasingly composed of smaller datasets to leverage the expertise of specific subteams, greater collaboration across groups is necessary. This can lead to increased friction in adopting new tools and exploration patterns (Kandel et al., 2012), as stakeholders and collaborators must transition to new tooling simultaneously, or migrate in a manner that preserves data sharing capabilities.
@@ 5. Discussion
The reason why practitioners have not aligned on alternative tooling is not obvious. Practitioners across all domains recognize that there is a gap in the workflow:
@@
“Right now, if you want to curate high-quality data, you go through [each point] manually as an expert, which is not scalable [for] thousands of examples.” —T2
Practitioners are aware of and have tried the existing tools in this space. They are aligned on the properties that they want out of this tool (Table 2, Challenges), and these requests are being communicated to tooling teams:
“The kinds of requests we tend to get nowadays are about larger-scale dataset management, like mixture building. When you have a big selection, reviewing 10,000 rows is not what you want to do …That is much more amenable to summary review.” —T1
@@
5.0.1. The toolmakers’ hypothesis: the world is new.
When tool developers (T1-T4) described exploration workflows, they explained that there was a lack of alignment because the field is new:
@@
Second, our participants described a landscape where there was a lack of alignment (Doshi-Velez and Kim, 2017; Gilpin et al., 2018) across multiple topics such as objectives, metrics, and benchmarks, suggesting that the field and its principles are still emerging. The following are representative quotes from participants:
@@
T1, on LLM prompts: “There’s so many competing definitions of prompt quality…it’s a research north star that happens to be a major product priority. How can we improve this extremely important data set?’
M1, on training data: “The quality of data is subjective; a lot of people disagree…one person thinks it’s really high-quality data, but there’s no objective.”
T3, on evaluation data: “There’s not a framework for evaluating [data].. in a perfect world, there is well-articulated behavior (tone, subject matter, objective results)..”
@@
M1: “[Consider] search rankings…what makes for a good benchmark, how do we come to an agreement?”
E1: “If you’re doing simple classification, it’s easy to measure accuracy or precision or recall. But with generative models, evaluation is very subjective. Even the output of the model is subjective, so then, what’s going into the model- it’s really hard to say, is this better or worse?”
@@
T2: “Think about safety data curation…people can’t agree on criteria, let alone apply that criteria at scale.”
@@
T3: “What [data practitioners are] actually doing and what they communicate that they need are two very different things. What are they actually trying to do?”
5.0.2. The model developers’ hypothesis: there’s no tool that works for my use case.
Modeling and evaluation practitioners speculated that alignment was unlikely due to custom needs and requirements (Section 3.1).
“I think why [a spreadsheet is] so universal is that it’s so basic.. you can customize it to give this affordance that other tools may not give you.. it’s simple.” —E1
“We have tried so many [tools]. These tools are limiting is because they offer you exploration on only one aspect of [the data]…For me, they’re too specific.” —M2
Interestingly, when asked about the custom requirements for their use cases, practitioners listed similar requirements, which suggest that there may be opportunities for shared methods and evaluation frameworks. Some of these requirements include:
@@
It is likely the case that both the toolmakers’ and model developers’ hypotheses are true to some extent. There may be select opportunities for alignment as the field matures, and there are likely other problems that will require custom solutions. For example, there are specific tools being developed to address challenges that persist across datasets, such as safety and toxicity classification (Bellamy et al., 2018).
6. Conclusions and Future Work